Displaying report 1-1 of 1.
Reports until 18:29, Saturday 03 February 2024
Controls General (GQuEST)
Jeffrey Wack - posted 18:29, Saturday 03 February 2024 - last comment - 10:51, Friday 16 February 2024(11450)
piezo-controlled filter cavity open loop transfer function measurement

Attached are descriptions of the measurement setup used in log post 11449, as well as transfer function and noise data.

The piezo used here was Thorlabs, as described in 11449. The piezo transfer functions taken in log post 11373 were of the Noliac, would it be useful to take the same measurement of the Thorlabs piezo? Should it give us the same transfer function we will get by quotienting the open loop transfer function by the controller?

The noise measurements were taken with the laser locked, with the fan off and fan on.

 

 

Images attached to this report
Non-image files attached to this report
Comments related to this report
Jeffrey Wack - 10:51, Friday 16 February 2024 (11467)
Images attached to this comment
Displaying report 1-1 of 1.